OpenStudy (anonymous):

the area of a square of length(x+4)cm is equal to the area of a rectangle whose length is(2x+2)cm and width is xcm. Find x.

OpenStudy (noelgreco):

Can you draw a diagram of each shape?

OpenStudy (noelgreco):

Labeling the sides, of course.

OpenStudy (zzr0ck3r):

area of square = length^2 area of rec = length * width so (x+4)^2 = x(2x+2) x^2 +8x + 16 = 2x^2 + 2x -x^2 + 6x + 16= 0 x^2-6x-16 = 0 x^2-2x+8x-16 = 0 x(x-2) +8(x-2) = 0 (x-2)(x+8) = 0 x-2 = 0 x+8 = 0 x = 2 or -8 but -8 makes no sence in the context so x = 2

OpenStudy (zzr0ck3r):

wow just noticed this was 3 days old...strange.

OpenStudy (zzr0ck3r):

@loveth2093 why are you still bumping this? does it not make sense?